Commuting students. At a well-known university, 1/4 of the undergraduate students commute, and 1/3 of the graduate students commute. One-tenth of the undergraduate students drive more than 40 miles daily, and 1/6 of the graduate
students drive more than 40 miles daily. If there are twice
as many undergraduate students as there are graduate
students, then what fraction of the commuters drive more
than 40 miles daily?
: Commuting students. At a well-known university, 1/4 of the undergraduate students commute, and 1/3 of the graduate students commute. One-tenth of the undergraduate students drive more than 40 miles daily, and 1/6 of the graduate
students drive more than 40 miles daily. If there are twice
as many undergraduate students as there are graduate
students, then what fraction of the commuters drive more
than 40 miles daily?

Answer by ankor@dixie-net.com(4684)

Commuting students. At a well-known university, 1/4 of the undergraduates commute; 1/3 of the graduates commute. 1/10 of the undergraduates drive more than 40 miles daily and 1/6 of the graduates drive more than 40 miles daily. If there are twice as many undergrads as there are grad students, then what fraction of the commuters drive more than 40 miles daily?
:
Let's use a convenient number for the total students: 720 has a lot of factors
:
It says,"there are twice as many undergrads as there are grad students,"
That would give us us 480 undergrads, and 240 grads
:
1/4 of the undergraduates commute;
1/4 * 480 = 120
:
1/3 of the graduates commute.
1/3 * 240 = 80
:
1/10 of the undergraduates drive more than 40 miles daily
1/10 * 480 = 48
:
1/6 of the graduates drive more than 40 miles daily.
1/6 * 240 = 40
:
what fraction of the commuters drive more than 40 miles daily?
Total commuters: 120 + 80 = 200
Total commuters drive more than 40 mi: 48 + 40 = 88
:
Then: 88/200 = 11/25 of the commuters commute over 40 mi